OpenStudy (anonymous):

Can somebody help me find the axis of symmetry for g(x) = 5x2 − 10x + 7

OpenStudy (anonymous):

x=1 it comes from -b/2a in general form ax^2+bx+c

OpenStudy (anonymous):

-(-10)/2(5)=10/10 10/10=1?

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

yes thats correct, no problem!

OpenStudy (anonymous):

Could you help me with another?

OpenStudy (anonymous):

I'll start a new question